OpenStudy (anonymous):

Can anyone help me with this problem? PLEASE HELP! 1. Given the function f(x) = 4x ‐ 8 a. Find the rate of change between the two stated values for x: 5 to 7 b. Find the equation of a secant line containing the given points:(5,f(5)) and (7,f(7)) 2. Graph the following function using transformations. Be sure to graph all of the stages on one graph. State the domain and range. For example, if you were asked to graph y= x^2+1 using transformations, you would show the graph of y = x^2 and the graph shifted up 1 unit. Please do not show only the final graph. y= -√ x-6

OpenStudy (anonymous):

it is a line, so the secant line is the same as the original line

OpenStudy (anonymous):

the rate of change is the same as the slope of the line, which is 4

OpenStudy (anonymous):

\[f(5)=4\times 5-8=12\] \[f(8)=4\times 8-8=24\] the rate of change is \[\frac{24-12}{8-5}=\frac{12}{3}=4\]

OpenStudy (anonymous):

How do I graph this? Or am I supposed to?